OpenStudy (anonymous):

Need help remembering how I solve for theta for arctan(theta) = (cos(phi))/2

OpenStudy (anonymous):

\[\arctan(\theta) = \frac{ \cos(\phi) }{ 2 }\] solve for theta

OpenStudy (anonymous):

take the tangent of both sides? \[\arctan(\theta) = \frac{\cos(\phi)}{2}\]\[\tan(\arctan(\theta)) = \tan\left(\frac{\cos(\phi)}{2}\right)\]\[\theta=\tan \left( \frac{ \cos \phi }{ 2 } \right)\]

OpenStudy (anonymous):

dont think that's it. because inverse of tangent is cotangent

OpenStudy (anonymous):

that's the same thing I did.

OpenStudy (anonymous):

That's it. Cotangent is the reciprocal of tangent, not the inverse. Arctan is defined such that \[\arctan(\tan(x)) = x\] Consequently: \[\tan(\arctan(x)) = x\]

OpenStudy (anonymous):

ok i appreciate it.